(t - 4)(t + 3)
need some help solving this example. please show work

Answer by user_dude2008(1862)

(t - 4)(t + 3) = t*t+3t-4t-4*3=t^2+3t-4t-12=t^2-t-12

Answer: (t - 4)(t + 3)=t^2-t-12